s(x)=x3+x2+x+1 for 0≤x≤1, s(x)=-x3+7x2-5x+3 for 1<x≤2. Is the function a cubic spline?
Answer #1

solution: Given that scx) = x² + x²+x+1 Sex) = -x + 1x² 5x+3 osaal <AS2 (0,1) (1,2) Now we can write, on (0,1) S(X) = (So(x)